OpenStudy (anonymous):

Use implicit differentiation to find the slope of the tangent line to the curve xy^3+xy= 20 at the point ( 10 , 1 ).

OpenStudy (amistre64):

what have you done so far?

OpenStudy (anonymous):

\[\frac{ d }{ dx }\left( x^ay^b \right)=x^a \left( by ^{b-1}\frac{ dy }{ dx } \right)+y^b ax ^{a-1}\]